# MapGPT: Map-Guided Prompting for Unified Vision-and-Language Navigation

### 14 Jan 2024 v1

@@ 3.2 Map-Guided Prompting
For the VLN task, previous work [49, 12, 3] has demonstrated the effectiveness of online constructed maps for navigation. However, how to construct the maps and transform them into a certain form to prompt LLMs has not been investigated in this domain. In fact, there is also limited research exploring the capability of LLM as an agent in understanding a real-world environment in the field of LLM. In this paper, we propose a novel map-guided prompting approach that encourages GPT agents to engage in global exploration of the environment, as shown in Figure 2(b).
@@
In the setting of the VLN task, the agent has never explored the entire environment and must construct a map based on its own observations online. We store the map as a dynamically updated graph, following the graph-based method DUET [12]. At each step $t$, we record all observed nodes along the navigation trajectory and their connectivity into the graph $G_{t}=\{V_{t},E_{t}\}$, where $V_{t}=\{v_{t,i}\}_{i=1}^{K}$ are a series of $K$ observed nodes marked with the index $i$ in the order of observations. All the edges between these observed nodes are recorded in $E_{t}$. At any step $t$, given the current location, the simulator will provide several neighboring nodes that are currently navigable. These new nodes and edges will be utilized for updating the graph from $G_{t-1}$ to $G_{t}$.
3.2.1 Converting Map into Prompt
@@
“Trajectory: Place {$en_{0}$} {$en_{1}$}{$en_{t}$},
where each $en_{j}$ corresponds to a node $v_{t,i}\in V_{t}$ stored in the order of observation. Thus, we consider $i$ as the ID of this place and fill it into the template to denote $en_{j}$.
@@
We transform the connectivity between places into the form of prompts as well, aiding the agent in understanding the spatial structure of the scene. Since the connectivity can only be observed at explored nodes, we always start with “Place {$en_{t}$} is connected with …”. All IDs corresponding to the neighboring accessible nodes of these exploded nodes will be listed using the following template:
@@
Place {$en_{0}$} is connected with Places {$an_{0}^{0}$}, {$an_{0}^{1}$}, …
Place {$en_{1}$} is connected with Places {$an_{1}^{0}$}, {$an_{1}^{1}$}, …
Place {$en_{t}$} is connected with Places {$an_{t}^{0}$}, {$an_{t}^{1}$}, …”,
where all the nodes should be filled with the place ID. Note that this map connectivity does not need to be updated if the agent decides to backtrack and revisit some previously explored nodes.
@@
The final step involves adding descriptions to each node of this topological map, enabling the agent to refer to them for path planning. As we have already provided descriptions of currently accessible nodes in the action space, and the selected actions are also included in the history to form explored nodes, there is no need to repeat these descriptions. It is sufficient to simply add the location IDs in the action space at each step. Specifically, each $a_{t,i}$ in the action space $A_{t}$ is reformulated as
{label} {directional_phrase} Place {$an_{t}^{i}$}: {$o_{t,i}$.
The agent can therefore find the corresponding descriptions of explored nodes and accessible nodes in history $H_{t}$ and action space $A_{t}$, respectively. However, we still have some unexplored and currently inaccessible nodes that are important, especially when the agent encounters obstacles in exploration and needs to revisit previous nodes for re-exploration. It is essential to record descriptions of these inaccessible nodes, as they provide supplementary information to assist the agent in backtracking to the most suitable node. We record these descriptions using the following template:
“Supplementary Info:
Place {$un_{0}$}: <{scene_0}>
Place {$un_{1}$}: <{scene_1}> …”,
where we retain only the scene descriptions for the agent’s reference. This is in line with the idea of coarse-scale representation in DUET [12]. Additionally, if there are no such nodes present in the map currently, we will directly set the prompt as “Supplementary Info: Nothing yet”.
3.3 Multi-Step Path Planning
@@
Inspired by the above insight and benefiting from the utilization of maps, we directly devise a planning module that demands the agent to explicitly generate its own multi-step path planning and dynamically update it at each step. Concretely, the agent is required to combine the current observation and thought to update a new multi-step path planning. The entire process is iterative, where the planning output of the current step serves as the input to the next step, allowing the agent to refer to previous plans. Therefore, at step $t$, the agent should refer to the last planning $P_{t-1}$, which is formulated as a prompt:
“Previous Planning: {$P_{t-1}$},
where $P_{0}$ is set as “Navigation has just started, with no planning yet”.
We have simultaneously modified part of the “output requirements” in the task description $D$ as follows: “Based on your ‘Map’, ‘Previous Planning’ and current ‘Thought’, you also need to update your new multi-step path planning to ‘New Planning’ ”. Therefore, we can parse the new planning $P_{t}$ in each step of the LLM output. In summary, the proposed MapGPT that combines map $M_{t}$ and path planning $P_{t-1}$ can be defined as follows:
$\small T_{t},P_{t},a_{t}=LLM(PM(D,I,H_{t},O_{t},A_{t},M_{t},P_{t-1})).$ (2)
In addition to the widely-used thought and action on the previous agents, MapGPT outputs additional multi-step planning information. This can help the agent perform path planning more effectively and strive to execute or update it automatically. A detailed analysis of these capabilities is presented in Section 4.4.
@@
NavGPT [48] GPT-3.5 2 w/ 8.02 26.4 16.7 13.0
MapGPT (Ours) GPT-3.5 1 w/o 8.48 29.6 19.4 11.6
DiscussNav [24] GPT-4 5 w/o 6.30 51.0 37.5 33.3
MapGPT (Ours) GPT-4 1 w/o 5.80 61.6 41.2 25.4
MapGPT (Ours) GPT-4V 1 w/o 5.84 56.0 44.9 35.9
Table 1: Results on 72 various scenes of the R2R dataset. “Exp#” refers to the number of GPT experts and “Dist” refers to the utilization of distance information.
@@ 4.1 Experimental Settings
We choose two datasets, R2R [5] and REVERIE [32], to validate our MapGPT since they have distinct instruction styles. R2R provides detailed step-by-step instructions while REVERIE only offers a high-level description of finding the target object, which usually requires more exploration in the environment. To unify the prompt system, we focus only on navigation performance, which involves finding the correct location or object to stop, while neglecting the object grounding sub-task in REVERIE. We therefore adopt several evaluation metrics for navigation, including Navigation Error (NE, the distance between agent’s final location and the target location), Success Rate (SR), Oracle Success Rate (OSR, SR given Oracle stop policy), and SR penalized by Path Length (SPL).
Implementation Details
Most of the experimental results of MapGPT are achieved using GPT-4 [27]. For visual observations, we follow NavGPT [48] and utilize BLIP-2 [23] to provide a caption for the observation, and employ Faster R-CNN [35] pretrained on the Visual Genome dataset [22, 4] to detect existing objects. To transfer our MapGPT to the REVERIE dataset, we only make some simple yet necessary modifications, which demonstrates that our MapGPT is unified in the VLN task. Additionally, as shown in Table 1, MapGPT can be applied to multiple LLMs and has built the first GPT-4V-based [28, 29] agent in the VLN field. More details can be found in Appendix A.
@@
Training Seq2Seq [5] 7.81 28 21 -
Speaker Follower [16] 6.62 45 35 -
EnvDrop [39] 5.22 - 52 48
Pretraining PREVALENT [19] 4.71 - 58 53
Rec-VLN-BERT [20] 3.93 69 63 57
HAMT [11] 2.29 73 66 61
DUET [12] 3.31 81 72 60
BEVBert [3] 2.81 84 75 64
Zero-shot DUET (Init. LXMERT [38]) 9.74 7 1 0
NavGPT [48] 6.46 42 34 29
MapGPT (Ours) 6.29 57.6 38.8 25.8
Table 2: Results on the validation unseen set of the R2R dataset under the single-prediction setting. MapGPT outperforms two non-pretrained methods and the zero-shot NavGPT.
4.2 Experimental Results
Comparison on the R2R Dataset
As shown in Table 1, we follow the settings of NavGPT [48] and DiscussNav [24] to first evaluate the performance of our proposed MapGPT compared to these zero-shot agents on 72 scenes of the R2R dataset. Except for 11 scenes from the validation unseen set of R2R, we also adopt 61 scenes from the R2R training set for the zero-shot setting. We randomly sample one path for each scene, with three instructions per path, resulting in a total of 216 instructions. By following this setting, we can evaluate the performance of these zero-shot agents across various scenarios. In addition, DiscussNav incorporates a general prompt technique to enhance performance. Specifically, it repeats the generation process of GPT-4 for five times and then leverage an additional expert to fuse these output results. However, this undoubtedly increases both the time and API costs significantly. Therefore, to ensure a fair comparison, we only evaluate the performance of our MapGPT, and previous NavGPT and DiscussNav under the single-prediction setting.
Through prompting GPT-4 to perform the VLN task, our MapGPT with only one GPT expert, surpasses DiscussNav in three metrics, decreasing navigation error (NE) by 0.5, gaining a 3.7% improvement in SR, and even achieving a boost of over 10% in OSR. MapGPT (GPT-4 based) has limited performance on the SPL metric, which could be attributed to two factors. Firstly, converting visual observations into textual descriptions results in information loss, causing the agent to struggle in accurately finding the target to stop. Secondly, map-guided prompts encourage the agent to continue global exploration when encountering uncertain situations, which increases the length of the navigation path, thus impacting the SPL metric. On the other hand, the MapGPT based on GPT-4V avoids the issue of information loss and achieves a performance of 35.9% on SPL. We also feed our proposed map-guided prompts into GPT-3.5, and it outperforms NavGPT’s results based on GPT-3.5. These experiments can prove that our proposed map-guided prompting is unified across different LLMs.
@@ 4.2 Experimental Results
Training Seq2Seq [5] w/o 8.07 4.20 2.84
RCM [44] w/o 14.2 9.29 6.97
SMNA [25] w/o 11.3 8.15 6.44
FAST-MATTN [32] w/o 28.2 14.4 7.19
Pretraining HAMT [11] w/o 35.4 31.6 29.6
DUET [12] w/ 50.0 45.8 35.3
Zero-shot MapGPT (Ours) w/o 42.6 28.4 14.5
Table 3: Comparisons on a randomly sampled subset from the validation unseen set of the REVERIE dataset. Note that we have retested the released HAMT and DUET on this same subset.
We further compare the navigation performance between the proposed MapGPT and previous NavGPT on a larger validation unseen set with 11 scenes and 783 trajectories. As shown in Table 2, due to data distribution shift, the success rate of MapGPT is slightly lower (38.8%) compared to the results on the 72 scenes. Compared to NavGPT (GPT-4 based), MapGPT (GPT-4 based) exhibits a noticeable reduction in NE, leading by 4.8% in SR, and a substantial 15.6% improvement in OSR. This may be attributed to our proposed map-guided prompt, which has raised the upper limit of agent navigation ability (OSR). However, due to the need for a fair comparison, we utilize the same vision models as NavGPT, which constrains the final success rate (SR) of the agent in identifying the targets.
Moreover, this standard evaluation set can also be utilized for comparison with previous learning-based agents. In Table 2, we categorize previous methods into two types, namely training on the R2R dataset only and fine-tuning a pretrained model on the training set of the R2R dataset. As a zero-shot approach without utilizing any VLN data for training, MapGPT has demonstrated significant advancements over two learning-based methods, showcasing its tremendous potential.
In summary, our proposed MapGPT is highly effective, as it can be applied simultaneously to three GPT models, achieving the best performance in a zero-shot setting. Besides, it only utilizes a single GPT navigation expert and does not include additional distance information.
@@
$\checkmark$ $\times$ $\times$ $\times$ 6.62 50.0 27.8 15.3
$\checkmark$ $\checkmark$ $\times$ $\times$ 6.49 49.5 32.9 19.4
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 6.40 59.7 37.5 24.8
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 5.80 61.6 41.2 25.4
Table 4: The performance impacts of different visual information extraction methods, incorporating map-guided prompts, and performing explicit multi-step path planning.
Figure 3: An example on the REVERIE dataset showcases MapGPT’s various abilities, including map understanding and global exploration (highlighted in yellow), multi-step path planning and the updating mechanism (comparing green and blue highlights, MapGPT plans to explore places 8, 9, and 10 at step 2, and gradually updates the progress until step 4), as well as backtracking capability (place 10 is not accessible from 9, thus backtracking to place 1 first).
Comparison on the REVERIE Dataset
Benefiting from the flexible single-expert system and the utilization of a universal map-guided prompt, we can effortlessly transfer MapGPT to the REVERIE dataset, adapting to different instruction styles. The validation unseen set of the REVERIE dataset contains 3,521 instructions in total. Considering the API costs and easier comparison for future work, we randomly sample 500 instructions from the validation unseen set for the zero-shot setting of the REVERIE benchmark. Besides, we have retested the released HAMT and DUET models on these 500 instructions for a fair comparison.
As shown in Table 3, the zero-shot MapGPT (GPT-4 based) exhibits greater competitiveness on the REVERIE, significantly outperforming some training-only models across all metrics. Moreover, when compared to HAMT which benefits from the pretraining and fine-tuning process, MapGPT demonstrates a highly competitive performance as well, surpassing by 7.2% in the OSR metric, though experiencing a 3.2% drop in SR. MapGPT’s inferior performance compared to DUET could potentially be attributed to the utilization of additional precise GPS coordinate information in DUET, which is not adopted in our MapGPT.

### 25 Feb 2024 v2

@@ 3.2 Map-Guided Prompting
For the VLN task, previous work Zhu et al. (2021); Chen et al. (2022); An et al. (2023) has demonstrated the effectiveness of online constructed maps for global navigation. However, how to construct the maps and transform them into a certain form to prompt LLMs has not been investigated in this domain. Additionally, we observe that GPT-4V struggles slightly in its attempts to understand navigation environments based on multiple precise coordinates. Therefore, we propose a novel map-guided prompting approach that converts topological relationships of a map into textual prompts, as shown in Figure 2(b).
@@
In the VLN task, the agent has never explored the entire environment and must construct a map based on its own observations online. We store the map as a dynamically updated graph, following the graph-based method DUET Chen et al. (2022). At each step $t$, we record all observed nodes along the navigation trajectory and their connectivity into the graph $G_{t}=\{V_{t},E_{t}\}$, where $V_{t}=\{v_{t,i}\}_{i=1}^{K}$ are a series of $K$ observed nodes marked with the index $i$ in the order of observations. All the edges between these observed nodes are recorded in $E_{t}$. At any step $t$, given the current location, the simulator will provide several neighboring nodes that are currently navigable. These new nodes and edges will be utilized for updating the graph from $G_{t-1}$ to $G_{t}$.
3.2.1 Constructing Maps with Prompts
@@
“Trajectory: Place {$en_{0}$}{$en_{t}$},
where each $en_{j}$ corresponds to a node $v_{t,i}\in V_{t}$ stored in the order of observation. Thus, we consider the index $i$ as the ID of the node and fill it into the template to denote $en_{j}$.
@@
Unlike DUET, which also converts precise GPS coordinates into embeddings for graph learning, we only retain the topological relationships of the map nodes since we discover that it seems challenging for GPT to understand precise coordinate data. These topological relationships are transformed into textual prompts, making it easier to comprehend spatial structures. Since the connectivity can only be observed at explored nodes, we always start with “Place {$en_{t}$} is connected with …”. All IDs corresponding to the neighboring accessible nodes of these exploded nodes will be listed using the following template:
@@
Place {$en_{0}$} is connected with Places {$an_{0}^{0}$}, …
Place {$en_{1}$} is connected with Places {$an_{1}^{0}$}, …
Place {$en_{t}$} is connected with Places {$an_{t}^{0}$}, …”,
where all the nodes should be filled with their node IDs. Note that this map connectivity does not need to be updated if the agent decides to backtrack and revisit some previously explored nodes.
@@
The final step involves adding an annotation to each node of this topological map, enabling the agent to refer to them for path planning. As we have already provided currently accessible nodes in the action space, and the selected actions are also included in the history to form explored nodes, there is no need to repeat them. It is sufficient to simply add the node IDs in the action space at each step. Specifically, each $a_{t,i}$ in the action space $A_{t}$ is reformulated as
{label} {direction} Place {$an_{t}^{i}$}: {$o_{t,i}$.
The agent can therefore find the corresponding explored nodes and accessible nodes in history $H_{t}$ and action space $A_{t}$ respectively. However, we still have some unexplored and currently inaccessible nodes that are important, especially when the agent encounters obstacles in exploration and needs to revisit previous nodes for re-exploration. These inaccessible nodes are considered as supplementary information to assist the agent in backtracking to the most suitable node. In “Supplementary Info” prompts, we record their raw images for the one-stage system, whereas in the two-stage system, it can be replaced with the corresponding scene descriptions.
3.3 Adaptive Path Planning
@@
Inspired by the above insight and benefiting from the utilization of maps, we propose an adaptive planning module that demands the agent to dynamically generate and update its multi-step path planning at each step. Concretely, the agent is required to combine the thought, map, and previous planning to adaptively update a new multi-step path planning. The entire process is iterative, where the planning output of the current step serves as the input to the next step, allowing the agent to refer to previous plans. Therefore, at step $t$, the agent should refer to the last planning $P_{t-1}$, where $P_{0}$ is set as “Navigation has just started, with no planning yet”.
In summary, the proposed MapGPT that combines map $M_{t}$ and path planning $P_{t-1}$ can be defined as follows:
$\begin{split}T_{t},P_{t},a_{t}=&LLM(PM(D,I,H_{t},O_{t},A_{t},\\ &M_{t},P_{t-1})).\end{split}$ (2)
In addition to the widely-used thought and action on the previous agents, MapGPT outputs additional multi-step planning information. Thanks to the powerful reasoning abilities of LLMs, the agent can focus on multiple potential nodes or sub-goals during planning, instead of being limited to predicting only one optimal choice from the global action space in a probabilistic manner without interpretability, as in the previous DUET model. Furthermore, the agent can adaptively update its plan, choosing to continue exploring sub-goals or backtrack to a previous node for re-exploration, which enhances the agent’s navigation performance. Some analysis of these capabilities is presented in Section 4.
@@
NavGPT (Zhou et al.) GPT-3.5 3 8.02 26.4 16.7 13.0
MapGPT (Ours) GPT-3.5 1 8.48 29.6 19.4 11.6
DiscussNav (Long et al.) GPT-4 5 6.30 51.0 37.5 33.3
MapGPT (Ours) GPT-4 1 5.80 61.6 41.2 25.4
MapGPT (Ours) GPT-4V 1 5.62 57.9 47.7 38.1
Table 1: Results on 72 various scenes of the R2R dataset. “Exp” refers to the number of GPT experts.
@@ 4.1 Experimental Settings
We choose two datasets, R2R Anderson et al. (2018b) and REVERIE Qi et al. (2020b), to validate our MapGPT since they have distinct instruction styles. R2R provides detailed step-by-step instructions while REVERIE only offers a high-level description of finding the target object, which usually requires more exploration in the environment. To unify the prompt system, we focus only on navigation performance, which involves finding the correct location or object to stop, while neglecting the object grounding sub-task in REVERIE. We therefore adopt several evaluation metrics for navigation, including Navigation Error (NE, the distance between agent’s final location and the target location), Success Rate (SR), Oracle Success Rate (OSR, SR given Oracle stop policy), and SR penalized by Path Length (SPL).
4.2 Experimental Results
4.2.1 Results on the Room-to-Room Dataset
Various scenes. As shown in Table 1, We employ an identical sampled subset (72 scenarios, 216 trajectories) as in NavGPT’s experiment to evaluate the zero-shot performance across various scenarios. In addition, some of DiscussNav’s experiments fuse five repeated predictions to enhance performance. For a fair comparison, we evaluate the performance of our MapGPT, and previous NavGPT and DiscussNav under the single-prediction setting. For a two-stage system, MapGPT outperforms previous models when applied to different LLMs (including GPT-3.5 and GPT-4). When utilizing GPT-4V as a one-stage agent, combined with the proposed map-guided prompting with adaptive path planning, MapGPT achieves a success rate of 47.7%. MapGPT (GPT-4 based) has limited performance on the SPL metric, which could be attributed to the fact that map-guided prompting encourages the agent to continue global exploration when encountering insufficient textual descriptions. The agent has a longer navigation path, thus impacting the SPL metric. On the other hand, MapGPT based on GPT-4V avoids the issue of information loss during vision-to-text conversion and achieves an SPL of 38.1%.
@@ 4.2.1 Results on the Room-to-Room Dataset
Train Seq2Seq Anderson et al. (2018b) 7.81 28 21 -
Speaker Fried et al. (2018) 6.62 45 35 -
EnvDrop Tan et al. (2019) 5.22 - 52 48
Pretrain LangNav Pan et al. (2023) - - 43 -
PREVALENT Hao et al. (2020) 4.71 - 58 53
RecBERT Hong et al. (2021) 3.93 69 63 57
HAMT Chen et al. (2021b) 2.29 73 66 61
DUET Chen et al. (2022) 3.31 81 72 60
ScaleVLN Wang et al. (2023) 2.09 88 81 70
No NavGPT Zhou et al. (2023) 6.46 42 34 29
MapGPT (with GPT-4) 6.29 57.6 38.8 25.8
MapGPT (with GPT-4V) 5.63 57.6 43.7 34.8
Table 2: Results on the validation unseen set of the R2R dataset. MapGPT outperforms two non-pretrained methods and the zero-shot NavGPT.
Validation unseen set. We further compare the navigation performance between the proposed MapGPT and previous NavGPT on a larger validation unseen set with 11 scenes and 783 trajectories. As shown in Table 2, due to data distribution shift, the success rate of MapGPT is slightly lower (38.8% with GPT-4 and 43.7% with GPT-4V) compared to the results on the 72 scenes. Compared to NavGPT (GPT-4 based), MapGPT (GPT-4 based) exhibits a noticeable reduction in NE, leading by 4.8% in SR, and a substantial 15.6% improvement in OSR. Obviously, our proposed map-guided prompting with adaptive planning has raised the upper limit of agent navigation ability (OSR), thus increasing final success rate. However, due to information loss in the two-stage pipeline, MapGPT (GPT-4 based) tends to continue exploring when encountering insufficient description and has a weaker SPL performance. Nevertheless, the GPT-4V-based agent does not suffer from this issue, thus achieving 34.8% in SPL.
4.2.2 Results on the REVERIE Dataset
Benefiting from the flexible single-expert system and the utilization of a universal map-guided prompting and planning, we effortlessly apply MapGPT to the REVERIE dataset with different instruction styles. Considering the API costs and easier comparison for future work, we randomly sample 500 instructions from the validation unseen set for the zero-shot setting of the REVERIE benchmark.
Validation unseen set. As shown in Table 3, MapGPT exhibits greater competitiveness on REVERIE (31.6% SR), significantly outperforming zero-shot NavGPT and some training-only models across all metrics. Moreover, when compared to HAMT which benefits from the pretraining and fine-tuning process, MapGPT demonstrates a highly competitive performance as well.
Backtracking ratio. We also compare the backtracking ratio of NavGPT and our MapGPT on REVERIE. MapGPT experiences at least one instance of backtracking in 49% of cases, and among them, there is an 80% probability of successfully correcting its erroneous navigation path at least once. As a comparison, NavGPT, which does not utilize any maps and planning, performs poorly on this metric. Only 18% of cases involve backtracking to visited nodes, and among these cases, 53% of instances successfully correct the previous exploration error at least once.
@@ 4.2.2 Results on the REVERIE Dataset
Train Seq2Seq Anderson et al. (2018b) 8.07 4.20 2.84
RCM Wang et al. (2019) 14.2 9.29 6.97
SMNA Ma et al. (2019) 11.3 8.15 6.44
FAST-MATTN Qi et al. (2020b) 28.2 14.4 7.19
Pretrain HAMT Chen et al. (2021b) 35.4 31.6 29.6
DUET Chen et al. (2022) 50.0 45.8 35.3
No NavGPT Zhou et al. (2023) 28.3 19.2 14.6
MapGPT (with GPT-4) 42.6 28.4 14.5
MapGPT (with GPT-4V) 36.8 31.6 20.3
Table 3: Comparison on a randomly sampled subset from the validation unseen set of the REVERIE dataset. Note that we have retested the released HAMT and DUET on this same subset.
@@
GPT-4 $\times$ $\times$ 6.49 49.5 32.9 19.4
Topological $\times$ 6.40 59.7 37.5 24.8
Topological Adaptive 5.80 61.6 41.2 25.4
GPT-4V $\times$ $\times$ 5.96 58.8 42.6 34.7
Coordinate $\times$ 6.12 55.1 41.2 32.8
Topological $\times$ 5.89 56.5 44.9 36.5
Topological Action 5.82 58.3 45.4 35.6
Topological Adaptive 5.62 57.9 47.7 38.1
Table 4: Ablation of different map or planning designs on 72 various scenes of the R2R dataset.
Figure 3: A successful case on REVERIE showcases MapGPT’s (GPT-4V based) various abilities, including global exploration (blue), map understanding (yellow), and adaptive multi-step path planning (green). The six images on the right represent six unexplored places at step 4. Among these, MapGPT focuses on four possible places and systematically explores them until it discovers the bathroom when moving to place 8.